1-INVENTORY MANAGEMENT

Williams & Sons last year reported sales of $12 million, cost of goods sold (COGS) of $8 and an inventory turnover ratio of 2. The company is now adopting a new inventory system. If the new system is able to reduce the firm's inventory level and increase the firm's inventory turnover ratio to 4 while maintaining the same level of sales and COGS, how much cash will be freed up? Do not round intermediate calculations. Round your answer to the nearest dollar.

2-Receivables Investment

Medwig Corporation has a DSO of 34 days. The company averages $10,000 in sales each day (all customers take credit). What is the company's average accounts receivable? Round your answer to the nearest dollar.
Answer #1

Answer 1

Sales =$12 million

COGS =$8 million

Inventory Turnover Ratio = 2

Hence Inventory Level = COGS/ Inventory Turnover = 8/2=$4 million

As per new inventory system, Inventory Level = COGS/ Inventory Turnover = 8/4=$2 million

Hence the New inventory is reduced from $4 million to $2 million. Hence cash freed up = 4-2=$2 million

Answer 2

DSO =34 days

Sales =$10000 each day

DSO = Accounts receivable *No. of days/ Sales (Since Sales is provided in per day, no of days =1)

Accounts Receivable = DSO* Sales = 34*10000 = $340,000